OpenStudy (anonymous):

inscribed in an equalateral riangle is a circle. the height of the circle is 12 times the square root of 3. what is the circumference?

OpenStudy (anonymous):

okay, let's write the given:\[h = 12\sqrt{?}\] the height is actually the diameter of the circle so : \[D = h = 12\sqrt{3}\] So you can find R : \[R = D/2 = 12\sqrt{3}/2 = 6\sqrt{3}\] substitute R in the general equation of the circumference which is: \[C = 2\pi r =12\sqrt{3}\pi\] Correct me if I'm wrong ^_^

OpenStudy (anonymous):

h= 12sqrt(3) * my bad :)

OpenStudy (anonymous):

the circumference of the circle is = \[2\prod_{}^{}r\] =2 * pi * (\[12\sqrt{3} / 2\]

OpenStudy (anonymous):

= 2 * pi * 6 sqrt(3)

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

correct answer ???

OpenStudy (anonymous):

did u get it?

OpenStudy (anonymous):

lol that's what I got, he hopefully got it now ^_^